OpenStudy (anonymous):

Help me once again please: 3+5i ----- 1-2i

OpenStudy (dumbcow):

what are we solving for?

OpenStudy (anonymous):

um question is "Evaluate the following"

OpenStudy (anonymous):

its that thing where i^1 = i i^2 = -1 i^3 = -i i^4 = 1

OpenStudy (dumbcow):

you could multiply by conjugate to get rid of i in the denominator \[\frac{(3+5i)(1+2i)}{(1-2i)(1+2i)}\]

OpenStudy (anonymous):

then expand? the problem and solve correct?

OpenStudy (dumbcow):

yes you should get (11i -7)/5

OpenStudy (anonymous):

i got the 11i-5 but where does the /5 come from?

OpenStudy (dumbcow):

when you multiply the bottom 1 -2i +2i -4i^2 1-4i^2 1-4(-1) 1+4 5

OpenStudy (anonymous):

Oh so multiplying both top and bottom doesn't cancel out the bottom. I see where i missed up, thank you very much :)